OpenStudy (anonymous):

a cattle train left miami and traveled toward new york. 14 hours later a diesel train left traveling at 45km/h in an effort to catch up to the cattle train. after traveling for four hours the diesel train finally caught up. what was the cattle train's average speed?

OpenStudy (anonymous):

45*4=180 180/18=10 10km/h

OpenStudy (anonymous):

C_train= ax+14a D_train=45x 4a+14a=180 solve for a